OpenStudy (anonymous):

What are the different types of methods used to solve a system of equations? Give an example of each method along with the solution.

jhonyy9 (jhonyy9):

- let substitution method 2x+y=3 y=-x from second equation you see that y=-x so than in this case you can substituting in the first equation in place of y the -x so than will get the value of x ,like 2x +(-x) = 3 2x-x=3 x=3 after this you know from second equation that y=-x so this mean that because you have got from first equation that x=3 so than y=-x will result y=-3 - hope so much that this is understandably sure right easy there is again the elimination method ,but you have said one described step by step with example